# FORGE: Force-Guided Exploration for Robust Contact-Rich Manipulation under Uncertainty

### v1

@@ III-C Early Termination
where $y_{t}$ is the true success label at time $t$. During training, episodes are always executed for the maximum length.
At deployment time, a confidence threshold, $p_{term}$, can be used to terminate the episode as soon as the policy believes it has succeeded: $a^{ET}_{t}>p_{term}$. This allows us to behave efficiently, a desirable property for industrial applications where it is important to reduce cycle times.
@@ IV-A Robot System
For all experiments, we use a Franka Panda robot and the FrankaPy [25] library for the impedance controller. All policies send targets to the controller at a rate of $15Hz$ while the controller operates at $1000Hz$. The Panda has joint-torque sensing, which is projected to EE-frame force values when needed by the policy: $\hat{F}^{ee}=J^{\dagger}(q)\tau^{ext}$, where $J^{\dagger}$ is the Jacobian pseudo-inverse and $\tau^{ext}$ are the estimated external joint torques [26]. Alternatively, a force-torque sensor can be used to estimate the contact force.
For the majority of our experiments, we calibrate the poses of each fixed object and artificially add noise. This allows us to analyze performance under known levels of pose estimation error. The calibration is done by guiding the arm to a successful pose for the respective task from which a nominal initial pose can be backed out. Unless otherwise reported, our real experiments use the same initial state randomization as in simulation (see Appendix -A).
For our last experiment, we assemble a planetary gear box (Section V-E) using the perception system from IndustReal [7] (retrained using data we collected). This model assumes the $z$-position of parts are known and uses a Mask-RCNN model [27] to estimate bounding boxes from which planar locations can be backed out. The perception errors in this system are largely caused by extrinsics calibration errors and minor bounding box prediction errors.
@@ IV-B Policy Training
Simulator: All policies are trained using the Factory simulation methods within IsaacGym [5]. In simulation, we estimate the external contact force experienced by the end-effector (akin to attaching a force-torque sensor on the robot, or projection from joint sensing of external torques as done on Franka robots). Noisy sensor values are used as policy input, whereas ground-truth sensor values are used to compute the excessive-force penalty. For all RL, we use recurrent PPO [28] with asymmetric actor-critic [29] to handle partial observability. Details on initial state and observation randomization can be found in Appendix -A.
Checkpoint Selection: For all tasks and models, we train three policies with separate random seeds. For the peg-insertion and gear-meshing tasks, all policies are deployed on the real-robot and reported results are averaged across the three policies. For the nut-threading task, we found that not all policies transferred reliably to the real world, even when high success rates are achieved in simulation. As such, for this task, we report results for the best of the three checkpoints (determined using $18$ test runs each).
Observation and Action Frames: To allow efficient generalization across the workspace, we assume actions and observations are defined relative to the tip of the fixed part. Specifically, the policy outputs a $4D$ relative transform from the tip of the fixed part (we assume upright parts). The control target is computed from the fixed part’s pose estimate and the relative pose from the policy. The policy output is bounded, which limits the operational volume of the end-effector (we allow targets to be up to $5cm$ away in all directions). Similar to the action space, all position observations are relative to the tip of the fixed part.
@@ V-A Baseline Comparisons
8mm Peg Success Rate $\uparrow$ Duration (s) $\downarrow$ $F_{mean}$ (N) $\downarrow$ $F_{max}$ (N) $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
FORGE 0.84 (0.05) 5.01 (0.17) 5.51 (0.24) 12.84 (0.37) 1.00 (0.0) 1.00 (0.0)
FORGE (No Force) 0.82 (0.06) 7.30 (0.42) 7.09 (0.35) 14.16 (0.39) 0.59 (0.08) 0.81 (0.07)
No FP (400kp) 0.64 (0.07) 6.06 (0.40) 6.94 (0.13) 11.94 (0.24) 0.83 (0.07) 0.92 (0.05)
No FP (600kp) 0.71 (0.07) 5.28 (0.35) 10.66 (0.15) 16.58 (0.32) 0.91 (0.05) 0.97 (0.03)
Baseline 0.64 (0.07) 5.09 (0.30) 11.81 (0.21) 17.93 (0.41) 0.97 (0.03) 1.00 (0.0)
Medium Gear Success Rate $\uparrow$ Duration (s) $\downarrow$ $F_{mean}$ (N) $\downarrow$ $F_{max}$ (N) $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
FORGE 0.98 (0.02) 6.34 (0.42) 7.95 (0.11) 15.10 (0.45) 0.95 (0.03) 1.00 (0.0)
FORGE (No Force) 0.93 (0.04) 9.02 (0.79) 8.49 (0.23) 14.68 (0.39) 0.60 (0.08) 1.00 (0.0)
No FP (400kp) 0.82 (0.06) 6.44 (0.23) 6.52 (0.14) 10.97 (0.24) 1.00 (0.0) 1.00 (0.0)
No FP (600kp) 0.73 (0.07) 6.99 (0.46) 9.48 (0.23) 15.73 (0.30) 0.94 (0.04) 0.97 (0.03)
Baseline 0.69 (0.07) 7.57 (0.50) 11.67 (0.45) 18.29 (0.40) 0.90 (0.05) 0.97 (0.03)
M16 Nut Success Rate $\uparrow$ Duration (s) $\downarrow$ $F_{mean}$ (N) $\downarrow$ $F_{max}$ (N) $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
FORGE 0.44 (0.07) 24.50 (1.35) 6.88 (0.14) 13.34 (0.17) 0.50 (0.11) 1.00 (0.00)
FORGE (No Force) 0.69 (0.07) 13.16 (0.66) 7.78 (0.17) 14.41 (0.28) 1.00 (0.0) 1.00 (0.0)
Baseline 0.20 (0.06) 27.89 (2.22) 7.32 (0.23) 16.73 (0.29) 0.11 (0.10) 1.00 (0.0)
TABLE I: Baseline Comparison FORGE (with and without force observations) is compared to baselines that do not include robust sim-to-real components. It is additionally compared to ablations that do not use an excessive-force penalty. Evaluations are performed over a total of $585$ trials on the real robot ($45$ per row). Standard errors are included in parentheses.
We first compare FORGE to a baseline method that does not include any FORGE components. Specifically, the Baseline method was not trained using force observations, an excessive-force penalty, or dynamics randomization; however, it was trained with the early termination procedure so meaningful episode durations could be reported. We also considered a version of FORGE that does not use force observations. The questions we seek to answer are: (Q1) Does FORGE lead to more robust sim-to-real transfer? (Q2) Does FORGE lead to policies with more desirable behavioural properties?
We considered one variation of each task from Section II-A: the $8mm$ Peg, the Medium Gear (with two abutting gears), and the M16 Nut. Along with policy success rate, used to measure robustness for Q1, the following metrics are reported to answer Q2:
  • Duration (s): The average trial length when using an early termination threshold of $p_{term}=0.9$.
@@
Each reported metric represents $45$ trials spread across $5$ workspace locations for the fixed part, and $3$ pose-estimation error levels ranging from $0-5mm$ (see Fig. 3). Similar randomization ranges were used as in simulation except for the in-hand part randomization where the part was placed centrally in the gripper. Results are reported in Table I.
One conclusion for Q1 is that FORGE outperformed the Baseline method for all tasks whether force is included or excluded from the observation space. This suggests that the primary benefits of FORGE come from the dynamics randomization and excessive-force penalty. For FORGE, comparison to the policy without force observation shows that although using force sensing was useful for the easier insertion and gear meshing tasks, it harmed performance for the nut-threading task. We hypothesize that this is because nut-threading policies rely more heavily on force observations in simulation; these policies would therefore be more sensitive to any sim-to-real gap for this observation modality.
Examining the behavioural metrics for Q2, we notice that FORGE used less force than the baseline and had minor improvements in trial duration. During experiments, we observed FORGE led to gentler interactions between the parts (see accompanying video). The reduced force produced by this policy was especially helpful for the M16 Nut which was more susceptible to slipping than the peg or gear.
V-B Noise Analysis
Figure 3: Perception Error For each task, we visualize what the different pose estimation errors look like overlaid on the fixed part.
Figure 4: Noise Analysis Performance broken down by level of pose error. Each subplot is a planar representation of the error levels where each ring corresponds to low (0-1mm), medium (1-2.5mm), and high (2.5-5mm) error. Success rate, stated in black text, is also represented by the shade of the corresponding ring. Dots represent x-y noise samples for successful (green) and failed (red) trials. FORGE results in good performance across tasks even with high error levels.
We next aim to answer (Q3): How is policy performance, in terms of success rate, affected by pose-estimation error? We use the same trials from the previous section, but show a breakdown of the results across different error levels. During each trial, artificial perception error was added to the fixed part’s pose (calibrated as described in Section IV-A [^3] Adding artificial noise allows us to better characterize performance across error level compared to a perception system whose bias and variance can be difficult to estimate and control.). A third of the trials fell in each of the three considered error levels (see Fig. 3): Low (0-1mm), Medium (1-2.5mm), and High (2.5-5mm). We considered $3D$ position error by sampling a perturbation vector with a radius uniformly sampled in the desired error range and a direction uniformly sampled from the unit-sphere.
Figure 4 visualizes the performance of the baseline policy vs. FORGE at different noise levels. Each subplot is a 2D representation of how much x-y error there was for each trial (z-dimension error not visualized). Each point corresponds to either a successful (green) or unsuccessful (red) trial. The color of the ring represents the success rate at the corresponding error levels (increasing outwards). For the M16 Nut task, we include results for the No Force ablation as this was the most robust policy that used dynamics randomization and a force threshold.
FORGE achieved high success rates ($>0.8$) for all tasks at low and medium error levels, even for the M16 nut. Although performance degraded with error $>2.5mm$, FORGE still significantly outperformed the baseline. With high error, the effects of contact are more pronounced because the robot may need to search longer before the task is complete.
V-C Force Analysis
Figure 5: Gains Analysis (90 trials, 8mm Peg) With force sensing (+Force), FORGE can achieve robust success rates (bottom) across varying controller gains at deployment time. Even with different gains, force sensing allows the policy to modulate its actions to achieve low contact forces (top).
Excessive-Force Penalty (Q4): In Table I, we compare to an ablation, No FP, that was trained without the excessive-force penalty of FORGE (but still used force observations and dynamics randomization). We used the same evaluation procedure as for FORGE but deployed with two different controller gains (we chose values at the lower and middle of the gain randomization range). Note that ablation results are not reported for nut-threading as we found that the nut always slipped out of the gripper. We found that policies deployed with the lower gains achieved similar average forces to FORGE while those deployed with higher gains naturally experienced more force. Both policies had lower success rates than FORGE which was deployed with controller gains at the middle of the randomization range.
Gains Robustness (Q5): To measure how robust FORGE is to controller gains, we performed an additional experiment where we varied the gains at deployment time and measured success rate. We compared FORGE and the No Force ablation to gain insight into how important force sensing is to limiting applied forces. The experiment was carried out for the $8mm$ peg task at a single workspace location, with medium pose estimation error and limited initial-state randomization. We considered $5$ proportional gain levels across the randomization range (corresponding to an $8N$ range in the maximum force the controller could apply) and each condition was evaluated $9$ times ($3$ runs per checkpoint).
In Fig. 5 (bottom), we see that FORGE achieves high success rates across a wide range of controller gains. However, performance is less consistent without force observations. In Fig. 5 (top), we use a box plot to show the spread of $F_{mean}$ across the $9$ trials of each condition. The dotted line shows the deployment force-threshold: $F_{th}=7.5N$. We see that when the force observation was included, contact force was consistently low across gains. However, without force observations, the spread of forces across episodes was high, often exceeding the threshold at higher gains. This highlights the importance of force sensing to enable the policy to effectively modulate the contact force.
V-D Success Prediction Analysis
Figure 6: Success Prediction Analysis Relationship between Delay Time and Success Rate for two early termination methods (generated by varying each method’s respective parameter: $T$ or $p_{term}$). The Pred Term method leads to lower delays than the Fixed Term method, especially at higher success rates. The vertical line shows a $0.8$ success rate.
To evaluate the early termination procedure, we ask: (Q6) How much efficiency is gained when the policy determines when to terminate? (Q7) Does success prediction, trained in simulation, transfer to the real world?
To answer Q6, we use Delay Time (s) to capture efficiency (lower values are better). Delay time measures the time between when success occurred and when the episode was terminated. We compare the proposed method (Pred Term) to a standard termination method that stops the policy after a fixed duration, $T$ (Fixed Term). Each method has a parameter that can be tuned to produce a different success rate (fraction of episodes that are successful when terminated). However, this will introduce a trade-off with delay time:
  • Fixed Term ($T$): Waiting too long is inefficient while terminating too early will harm success rates.
  • Pred Term ($p_{term}$): A high threshold can cause extra delay while a low threshold can affect success rate.
Fig. 6 is a simulated analysis that shows the relationship between Delay Time and Success Rate for each method. Each line was generated by measuring the success rate and corresponding delay time across a fine discretization of each method’s termination parameter. These were then sorted by success rate and plotted. [^4] Similar to an ROC plot, but higher areas above the curve are better. As a practitioner, one could choose a desired success rate and find the resulting delay.
Across all tasks, we see that the Fixed Term method leads to longer delays, especially at higher success rates (we plot a vertical line to show the $0.8$ success rate). The early termination action, $a^{ET}$, allows for dynamic episode lengths, leading to high success rates with smaller delay times.
@@ V-D Success Prediction Analysis
Finally to answer Q7, results in Table I show that early termination prediction transferred well to the real world. While the termination method tended to correctly identify successes for all models (high and often perfect recall), we see that precision was best when using force observations for the gear and peg tasks. This shows the benefit of force for sensing task completion: when the bottom of the socket has been reached, or the gear has been fully meshed.

### v2

@@ III-C Success Prediction
where $y_{t}$ is the true success label at time $t$. This reward can also encourage behaviours that elicit the underlying success state (e.g., pull upwards on the nut to check if it is threaded).
Early Termination: Efficient termination is a desirable property for industrial applications where cycle times matter. We want the policy to terminate as soon as the task has succeeded and no sooner. During training, episodes are executed for the maximum length. At deployment, a confidence threshold, $p_{term}$, can be used to terminate the episode: $a^{ET}_{t}>p_{term}$. See App. -D for analysis on the performance trade-offs for choosing $p_{term}$.
Force Threshold Tuning: For forceful insertion tasks, we may not know how much force is required. Consider snap-fit connectors which require deformation. The required force depends on material properties that may be unknown and could change with extended use. To tune the force threshold, we leverage success prediction. Conservatively, we start with a low threshold of $7.5N$, which helps avoid slip and damage. If policy execution reaches a timeout before success is predicted, we increase the threshold and try again. This can be done automatically, without manual resets, until success occurs.
@@ IV-A Robot System
We use a Franka Panda robot with the FrankaPy [29] library for impedance control. All policies send control targets at $15Hz$ while the controller operates at $1000Hz$. The Panda has joint-torque sensing, which is projected to EE-frame forces when needed by the policy [28]. Alternatively, a force-torque sensor could be used.
For the majority of our experiments, we calibrate the poses of each fixed object and artificially add noise. This allows us to analyze performance under known levels of position estimation error. The calibration is done by guiding the arm to a successful pose for the respective task from which a nominal initial pose can be backed out. Unless otherwise reported, our real experiments use the same initial state randomization as in simulation (see App. -A). For our last experiment, we assemble a planetary gear box (Sec. V-E) using the perception system from IndustReal [7] (see App. -F for more details).
@@ IV-B Policy Training
Simulator: All policies are trained using the Factory simulation methods within IsaacGym [5]. In simulation, we have access to external contact forces experienced by the end-effector (akin to what we have access to on the Panda). Noisy forces are used as policy input, whereas ground-truth forces are used to compute the excessive-force penalty. We use recurrent PPO [30] with asymmetric actor-critic [31] to handle partial observability. Details on initial state and observation randomization can be found in App. -A.
Checkpoint Selection: For all tasks and models, we train three policies with separate random seeds. On the real-robot, results are averaged across the three policies.
Observation and Action Frames: For generalization across the workspace, we assume actions and observations are relative to the fixed part. Specifically, the policy outputs a $4D$ relative transform from the tip of the fixed part (we assume upright parts). The control target is computed from the fixed part’s pose estimate and the relative pose from the policy. The policy output is bounded, limiting the operational volume of the end-effector (targets can be up to $5cm$ away in all directions). Similar to the action space, all position observations are relative to the tip of the fixed part.
IV-C Baselines and Ablations
We compare FORGE to two baselines:
IndustReal [7]: Policies trained using the IndustReal framework do not have velocity or force observations. A full description of the differences can be found in App. -C. The PLAI parameters from IndustReal are set to achieve similar maximum forces to what FORGE policies can command.
Baseline: Similar to FORGE but does not use force observations, dynamics randomization, or an excessive force penalty. However, it is trained with success prediction so that meaningful episode durations can be reported.
Ablations: In addition to the baselines, we also ablate each of the main components of FORGE for our sim-to-real analysis: Force (No Force), Dynamics Randomization (No DR), and Excessive Force Penalty (No FP). For the FORGE (No FP) model, which ablates the contact penalty reward term, we evaluate using two P-gain levels. Note that FORGE (No FP) results are not reported for nut threading as we found that the nut always slipped out of the gripper.
@@ V-A Baseline Comparisons
8mm Peg Success Rate $\uparrow$ Duration (s) $\downarrow$ $F_{mean}$ (N) $\downarrow$ $F_{max}$ (N) $\downarrow$ Precision $\uparrow$ Recall $\uparrow$ Delay (s) $\downarrow$
FORGE 0.84 (0.05) 2.82 (0.20) 5.51 (0.24) 12.84 (0.37) 1.00 (0.0) 1.00 (0.0) 2.19 (0.08)
FORGE (No Force) 0.82 (0.06) 3.18 (0.36) 7.09 (0.35) 14.16 (0.39) 0.59 (0.08) 0.81 (0.07) 4.12 (0.37)
FORGE (No DR) 0.91 (0.04) 2.03 (0.19) 6.28 (0.24) 13.08 (0.36) 1.00 (0.0) 0.98 (0.02) 2.51 (0.04)
FORGE (No FP, 400kp) 0.64 (0.07) 3.21 (0.35) 6.94 (0.13) 11.94 (0.24) 0.83 (0.07) 0.92 (0.05) 2.85 (0.40)
FORGE (No FP, 600kp) 0.71 (0.07) 2.88 (0.35) 10.66 (0.15) 16.58 (0.32) 0.91 (0.05) 0.97 (0.03) 2.40 (0.22)
Baseline 0.64 (0.07) 2.35 (0.27) 11.81 (0.21) 17.93 (0.41) 0.97 (0.03) 1.00 (0.0) 2.74 (0.06)
IndustReal 0.82 (0.06) 3.41 (0.24) 9.45 (0.14) 21.15 (0.26) N/A N/A 6.59 (0.24)
Medium Gear Success Rate $\uparrow$ Duration (s) $\downarrow$ $F_{mean}$ (N) $\downarrow$ $F_{max}$ (N) $\downarrow$ Precision $\uparrow$ Recall $\uparrow$ Delay (s) $\downarrow$
FORGE 0.98 (0.02) 3.14 (0.39) 7.95 (0.11) 15.10 (0.45) 0.95 (0.03) 1.00 (0.0) 3.20 (0.28)
FORGE (No Force) 0.93 (0.04) 3.06 (0.29) 8.49 (0.23) 14.68 (0.39) 0.60 (0.08) 1.00 (0.0) 5.96 (0.70)
FORGE (No DR) 0.87 (0.05) 3.42 (0.50) 7.15 (0.20) 13.94 (0.36) 0.90 (0.05) 1.00 (0.0) 3.04 (0.24)
FORGE (No FP, 400kp) 0.82 (0.06) 3.57 (0.26) 6.52 (0.14) 10.97 (0.24) 1.00 (0.0) 1.00 (0.0) 2.87 (0.16)
FORGE (No FP, 600kp) 0.73 (0.07) 3.08 (0.29) 9.48 (0.23) 15.73 (0.30) 0.94 (0.04) 0.97 (0.03) 3.91 (0.39)
Baseline 0.69 (0.07) 2.90 (0.37) 11.67 (0.45) 18.29 (0.40) 0.90 (0.05) 0.97 (0.03) 4.68 (0.41)
IndustReal 0.87 (0.05) 8.44 (0.61) 9.80 (0.16) 20.48 (0.26) N/A N/A 6.56 (0.61)
M16 Nut Success Rate $\uparrow$ Duration (s) $\downarrow$ $F_{mean}$ (N) $\downarrow$ $F_{max}$ (N) $\downarrow$ Precision $\uparrow$ Recall $\uparrow$ Delay (s) $\downarrow$
FORGE 0.69 (0.07) 11.38 (0.47) 7.82 (0.13) 14.52 (0.22) 0.74 (0.08) 0.74 (0.08) 6.54 (1.25)
FORGE (No Force) 0.40 (0.07) 14.09 (1.11) 8.34 (0.15) 15.04 (0.17) 0.33 (0.11) 0.33 (0.11) 11.48 (1.63)
FORGE (No DR) 0.56 (0.07) 11.12 (0.14) 7.65 (0.17) 14.10 (0.24) 0.72 (0.09) 0.86 (0.08) 8.10 (1.45)
Baseline 0.40 (0.07) 11.19 (0.24) 10.51 (0.51) 17.34 (0.36) 0.72 (0.11) 0.93 (0.07) 10.16 (1.69)
IndustReal 0.36 (0.07) 22.27 (0.64) 12.63 (0.31) 22.34 (0.33) N/A N/A 7.73 (0.64)
TABLE I: Baseline Comparison FORGE is compared to baselines that are not force-aware (IndustReal [7] and Baseline). It is also compared to ablations that do not observe force (No Force), a force penalty (No FP), or dynamics rand (No DR). Evaluations are performed over a total of $855$ trials on the real robot ($45$ per row). Standard errors are included in parentheses.
(Q1) Does FORGE lead to more robust sim-to-real transfer? (Q2) Do FORGE policies have more desirable behavioural properties?
Along with success rate, used to measure robustness for Q1, the following metrics are reported for Q2:
  • Duration (s): For successful episodes, time to reach a successful state (independent of success prediction).
@@
Each reported metric represents $45$ trials spread across $5$ workspace locations for the fixed part, and $3$ position-estimation error levels ranging from $0-5mm$ (see Fig. 3). Similar randomization ranges were used as in simulation except for the in-hand part randomization where the part was centered in the gripper. Results are reported in Table I.
One conclusion for Q1 is that FORGE outperformed the Baseline method for all tasks and IndustReal for both the gear meshing and nut threading tasks. Ablations show that that the primary performance gains of FORGE come from including force observations and the excessive force penalty. Although dynamics randomization did not significantly affect success rate, we later show it is important for robustness across controller gains.
Examining the behavioural metrics for Q2, we notice that FORGE used less force than both baselines and had significant improvements in trial durations when compared to IndustReal. During experiments, we observed FORGE led to gentler interactions between the parts (see accompanying video). The reduced force produced by this policy was especially helpful for the M16 Nut which was more susceptible to slipping than the peg or gear.
For FORGE, the main failure cases occurred when there was high position estimation error (above $1\sigma$ of the training noise, see next section). Parts got stuck on each other (peg insertion) or the nut was rotated before alignment with the bolt, causing the threads to miss. However, we found training unstable with noise above $\sigma=2.5mm$. Adopting a curriculum or adding additional sensing modalities (e.g., tactile sensors or wrist-cameras) may help address this.
V-B Noise Analysis
Figure 3: Perception Error For each task, we visualize what the different position estimation errors look like overlaid on the fixed part.
Figure 4: Noise Analysis Performance broken down by level of position error. Each subplot is a planar representation of the error levels where each ring corresponds to low (0-1mm), medium (1-2.5mm), and high (2.5-5mm) error. Success rate, stated in black text, is also represented by the shade of the corresponding ring. Dots represent x-y noise samples for successful (green) and failed (red) trials. FORGE results in good performance across tasks even with high error levels.
We next aim to answer (Q3): How is policy performance, in terms of success rate, affected by position-estimation error? We use the same trials from the previous section, but show a breakdown of the results across different error levels. During each trial, artificial perception error was added to the fixed part’s position (calibrated as described in Section IV-A [^2] Adding artificial noise allows us to better characterize performance across error level compared to a perception system whose bias and variance can be difficult to estimate and control.). A third of the trials fell in each of the three considered error levels (see Fig. 3): Low (0-1mm), Medium (1-2.5mm), and High (2.5-5mm). We considered $3D$ position error by sampling a perturbation vector with a radius uniformly sampled in the desired error range and a direction uniformly sampled from the unit-sphere.
Figure 4 visualizes the performance of IndustReal vs. FORGE policies at different noise levels. Each subplot is a 2D representation of how much x-y error there was for each trial (z-dimension error not visualized). Each point corresponds to either a successful (green) or unsuccessful (red) trial (only $xy$ corrindates of the error vector are visualized as dots). The color of the ring represents the success rate at the corresponding error levels (increasing outwards).
Although performance is comparable for the peg insertion task, FORGE outperformed IndustReal for the gear meshing and nut threading tasks at all noise levels. This demonstrates that force is a useful modality to robustly recover from larger amounts of position estimation error. Performance generally degraded with error $>2.5mm$ which is beyond $1\sigma$ of the observation noise added in simulation. With high error, the effects of contact are more pronounced because the robot may need to search longer before the task is complete.
V-C Force Analysis
Figure 5: Gains Analysis (180 trials, 8mm Peg) With force sensing, FORGE can achieve robust success rates (bottom) across varying controller gains at deployment time. Even with different gains, force sensing allows the policy to modulate its actions to achieve low contact forces (top).
Excessive-Force Penalty (Q4): In Table I, we compare to an ablation, No FP, that was trained without the excessive-force penalty of FORGE (but still used force observations and dynamics randomization). We used the same evaluation procedure as for FORGE but deployed with two different controller gains (we chose values at the lower and middle of the gain randomization range). We found that policies deployed with the lower gains achieved similar average forces to FORGE while those deployed with higher gains naturally experienced more force. Both policies had lower success rates than FORGE which was deployed with controller gains at the middle of the randomization range.
Gains Robustness (Q5): To measure how robust FORGE is to controller gains, we performed an additional experiment where we varied the gains at deployment time and measured success rate. We compare FORGE to IndustReal and multiple ablations. The experiment was carried out for the $8mm$ peg task at a single workspace location, with medium position estimation error and limited initial-state randomization. We considered $5$ proportional gain levels across the randomization range (corresponding to an $8N$ range in the maximum force the controller could apply) and each condition was evaluated $9$ times ($3$ runs per checkpoint).
In Fig. 5, we see that FORGE achieves high success rates while respecting the force threshold across a wide range of controller gains. However, performance is less consistent without force observations or dynamics randomization. In Fig. 5 (top), we use a box plot to show the spread of $F_{mean}$ across the $9$ trials of each condition. The dotted line shows the deployment force-threshold: $F_{th}=7.5N$. We see that when the force observation was included, contact force was consistently low across gains. However, without force observations, the spread of forces across episodes was high, often exceeding the threshold at higher gains. Similarly, the force exerted by IndustReal policies increased with controller gains. As IndustReal is not force-aware, achieving desired forceful properties requires tuning controller gains. Overall, these results highlight the importance of force sensing to enable the policy to effectively modulate the contact force.
V-D Success Prediction Analysis
To evaluate success prediction, we ask: (Q6) Does success prediction, trained in simulation, transfer to the real world? (Q7) Can success prediction be used to tune the force-threshold for tasks that require forceful insertion?
Sim-to-Real Transfer (Q6): To measure the effective of success prediction, we report additional metrics for each of the trials in Table I:
@@ V-D Success Prediction Analysis
  • Early Term. Delay (s): For successful episodes, how long after success occurred did the policy terminate.
@@
Results show that success prediction transferred well to the real world. The termination method correctly identified successes (high precision and recall) and worked best when using force observations for all tasks. We also see that delay times are shortest when using force observations. This shows the benefit of force for sensing task completion: when the gear has been fully meshed or the nut threads successfully engaged. Using success prediction also leads to shorter delays than IndustReal which uses a fixed duration.
Force Threshold Tuning (Q7): To evaluate the utility of success prediction for force threshold tuning, we introduced a new snap-fit task (see Fig. 6). In simulation, the snap-fit buckle was implemented with torsion springs for each of the clips. The stiffness of the springs was randomized to vary the amount of force needed for insertion. We also ensured the robot’s gains and force-threshold were randomized such that success was possible. In the real world, we used a snap-fit buckle that required $15N$ of force for insertion. We report real world results from running the automatic tuning procedure described in Sec. III-C. The initial force threshold was set to $7.5N$ and increased by $5N$ each policy execution until the policy predicted success. No initial state randomization or noise were added for these experiments.
Out of $10$ trials for the complete tuning procedure (each consisting of up to $3$ policy executions with increasing force thresholds), the tuning procedure succeeds $8$ times while successful insertion occurred $10$ times. The two failures were instances where the insertion succeeded, but the policy failed to predict success. Success occurred on the third execution $9/10$ times (it once occurred on the second trial), meaning the policy generally respected the force threshold (success should not occur until the force threshold exceeds $15N$). Furthermore, of the $29$ policy executions, the success prediction by the policy was correct $27$ times.
We also evaluated the task using a sufficiently high threshold and on a larger initial state distribution, similar to what was used during training in simulation. Here the success rate dropped to $6/10$, which we attribute to an unstable grasp leading to part slippage during contact. This reveals a limitation of having a single force threshold: for certain tasks, the optimal force threshold may vary depending on the phase of the task. For example, low forces are required until the buckle is aligned with the socket, only then is it safe to use high forces.
Figure 6: Simulated (left) and real (right) parts for a snap-fit insertion task which requires $15N$ of force to succeed. See the video for policy execution.